1. Compute for the interest of a loan of P250,000 dated January 1, 2016 and payable on November 13, 2016 assuming the rate is 12% using the following: (20 points)

a. Actual Time-Exact Interest Method

b. Approximate Time-Exact Interest Method

c. Actual Time-Ordinary Interest Method

d. Approximate Time-Ordinary Interest Method


1
Expert's answer
2021-02-23T12:38:49-0500

Principle=250,000

Rate=12%

Total number of days in 2016=366 days

Actual loan duration =317 days

Approximate loan duration =313 days

a)Actual Time-Exact Interest Method

Interest=250,000*0.12*317 days/366 days    

         =25,983.6065

b. Approximate Time-Exact Interest Method

Interest=250,000*0.12*313 days/366 days            

 =25,655.7377

c. Actual Time-Ordinary Interest Method

Interest=250,000*0.12*317 days/360 days             

 =26,416.6667

d. Approximate Time-Ordinary Interest Method

Interest=250,000*0.12*313 days/360 days            

  =26,083.3333